OpenStudy (anonymous):

I am having trouble I started math correspondance and I need a little help here is my first question (8xyto the power of 3)(7y to the power of -4)

OpenStudy (anonymous):

and...?

OpenStudy (anonymous):

I do not understand how to do this kind of math

OpenStudy (anonymous):

that is simply an expression, what do you want to do with it?

OpenStudy (anonymous):

I have to simplify the expressions

OpenStudy (anonymous):

you'll need a couple definitions: (a^b)(a^c) = a^(b + c) a^(-b) = 1/(a^b) a^(0) = 1